A bag has equal numbers of magenta, turquoise, and black marbles. Corey picked a
marble from the bag, replaced it, then picked another marble. Both times he did not select a magenta marble. If he repeats this process 250 times, predict how many repetitions will result in no magenta marbles?
 4
 83
 167
 250

Found 2 solutions by greenestamps, math_tutor2020:
Answer by greenestamps(13198)



P(magenta) = P(turquoise) = P(black) = 1/3

P(not magenta) = 2/3

P(not magenta twice in a row) = (2/3)(2/3) = 4/9

Approximate number of times this result is obtained when trying the experiment 250 times: 250(4/9) = 1000/9 = approximately 111

None of the answer choices is close to the right answer....


Answer by math_tutor2020(3816)


I agree with the answer by tutor greenestamps.

It's possible that your teacher mistakenly thinks the answer is 167 due to the computation 250*(2/3) = 166.667 which is approximate.
167 would be the answer if each trial consisted of picking one marble (instead of two).